### v5

# Optimization via a Control-Centric Framework

Abstract
Optimization plays a central role in intelligent systems and cyber-physical technologies, where speed and reliability of convergence directly impact performance. In control theory, optimization-centric methods are standard: controllers are designed by repeatedly solving optimization problems, as in linear quadratic regulation, $H_{\infty}$ control, and model predictive control. In contrast, this paper develops a control-centric framework for optimization itself, where algorithms are constructed directly from Lyapunov stability principles rather than being proposed first and analyzed afterward. A key element is the stationarity vector, which encodes first-order optimality conditions and enables Lyapunov-based convergence analysis. By pairing a Lyapunov function with a selectable decay law, we obtain continuous-time dynamics with guaranteed exponential, finite-time, fixed-time, or prescribed-time convergence. Within this framework, we introduce three feedback realizations of increasing restrictiveness: the Hessian-gradient, Newton, and gradient dynamics. Each realization shapes the decay of the stationarity vector to achieve the desired rate. These constructions unify unconstrained optimization, extend naturally to constrained problems via Lyapunov-consistent primal-dual dynamics, and broaden the results for minimax and generalized Nash equilibrium seeking problems beyond exponential stability. The framework provides systematic design tools for optimization algorithms in control and game-theoretic problems.
publicationid: pubid: 0000–0000/00\$00.00 © 202X IEEE
I Introduction
Optimization is a cornerstone of intelligent systems and cyber-physical technologies, where it underlies decision-making, resource allocation, and learning-based control. In modern control theory, optimization-based approaches are likewise central. Prominent examples include the linear quadratic regulator (LQR) [kalman_contributions_1960], $H_{2}$ and $H_{\infty}$ control [doyle_state-space_1989], and model predictive control (MPC) [rawlings_model_2024], where the control design is obtained by formulating and solving an optimization problem at each step of the process. These methods may be viewed as optimization-centric frameworks for control, since the optimization problem itself constitutes the foundation upon which control laws are built.
In contrast to optimization-centric approaches to control, this paper focuses on a control-centric framework for optimization itself, where continuous-time dynamics are designed as optimizers by directly invoking Lyapunov stability theory and its modern extensions. This perspective parallels the classical theory of control Lyapunov functions (CLFs), in which stabilizing feedback laws are constructed directly from the Lyapunov function. Foundational results include Artstein’s equivalence between CLF existence and feedback stabilizability [artstein_stabilization_1983], Sontag’s universal construction of continuous stabilizing feedbacks [sontag_lyapunov-like_1983, sontag_universal_1989], and Brockett’s necessary condition for continuous static stabilization [brockett_asymptotic_1983]. Building on this foundation, we introduce the notion of optimization Lyapunov functions (OLFs), which transfer these control-theoretic principles to the design of optimization algorithms. In particular, tools such as finite-time (FT) stability [bhat_finite-time_2000], fixed-time (FxT) stability [polyakov_nonlinear_2012], and prescribed-time (PT) stability [song_time-varying_2017] provide systematic ways to guarantee convergence with user-specified temporal properties. This control-centric viewpoint enables the construction of optimizer dynamics that meet explicit convergence requirements, rather than analyzing convergence only after an algorithm has been specified.
A substantial body of prior work has applied Lyapunov-based methods to study continuous-time optimization, but almost always in the form of analysis rather than design. Classical contributions include the pioneering work on saddle-point dynamics for constrained optimization in [arrow_studies_1958], and the primal-dual dynamics in [feijer_stability_2010]. The stability of saddle-point dynamics has since been analyzed under different convexity and monotonicity conditions [cherukuri_saddle-point_2017]. For unconstrained optimization, normalized and signed gradient dynamics with FT convergence were introduced by Cortés [cortes_finite-time_2006]. Subsequent extensions investigated projection-based dynamics for constrained convex problems [chen_convex_2018], as well as FT and FxT algorithms for more general problem classes [sanchez-torres_fixed_2014, li_fixed_2017, garg_fixed-time_2021]. For minimax problems and generalized Nash equilibrium (GNE) seeking, recent works have established FT or FxT convergence properties under appropriate assumptions [garg_fixed-time_2022, romero_finite-time_2020, you_unified_2024]. In parallel, Ross [ross_optimal_2019, ross_derivation_2023] showed how optimizer dynamics, including coordinate descent methods, can be systematically derived from optimal control principles, highlighting the long-standing interplay between control theory and optimization.
Despite this progress, most existing methods follow the pattern of proposing an optimization algorithm first and only then analyzing its convergence using control-theoretic tools. In contrast, our contribution is a framework for designing optimizers directly according to desired convergence properties. This perspective not only recovers known results as special cases but also closes several gaps in the literature. For example, while FT and FxT convergence results exist for unconstrained problems, comparable results for constrained optimization with PT guarantees remain largely unexplored. Our framework further enables the design of second-order continuous-time optimizers that exploit curvature information without requiring explicit Hessian inversion, thus extending beyond the current state-of-the-art.
The main contribution of this paper is the introduction of a control-centric framework that generalizes the classical notion of control Lyapunov functions (CLFs) to optimization through optimization Lyapunov functions (OLFs). The framework decouples the design of the stabilizing feedback from the specification of the decay law, allowing the convergence rate to be explicitly specified. The familiar exponential, FT, FxT, and PT behaviors are presented as representative examples of this general design capability. Within this formulation, we develop three continuous-time realizations: the Hessian-gradient, Newton, and gradient dynamics, which differ in structure and computational complexity but share the same Lyapunov-based synthesis principle. The Hessian-gradient dynamics incorporate second-order information through the Hessian matrix but avoid explicit matrix inversion by normalizing the gradient direction, providing a practical compromise between first- and second-order methods. The Newton dynamics utilize the full Hessian feedback to achieve fast convergence, whereas the gradient dynamics represent the simplest realization, based solely on first-order information. The framework is further extended to constrained programs, where Lyapunov-consistent Karush-Kuhn-Tucker (KKT) dynamics are constructed using smoothed stationarity mappings to guarantee convergence under the same selectable timing laws. Additionally, it is applied to minimax and GNE seeking problems, where analogous OLF designs yield unified stability guarantees that extend beyond asymptotic or exponential results. Finally, we briefly discuss how existing Lyapunov-preserving discretization schemes can maintain these continuous-time convergence properties in practice.
The remainder of this paper is organized as follows. Section II reviews the required background and introduces a family of convergence laws covering exponential, FT, FxT, and PT regimes. Section III formalizes the notion of OLFs and presents the control-centric methodology for constructing optimizer dynamics with selectable convergence rates. Sections IV-VI apply this methodology to different problem classes, beginning with constrained optimization and continuing with minimax formulations and GNE seeking problems. Section VII outlines how Lyapunov-preserving discretization schemes can retain the continuous-time stability properties of the proposed dynamics. Finally, Section VIII concludes the paper.
II Mathematical Background
In this section, we introduce the general mathematical framework and the assumptions that will be used throughout the paper. These assumptions are kept minimal and apply across the different problem classes treated in later sections. Additional assumptions required for specific cases (e.g., convexity, constraint qualifications) will be stated within their respective sections.
We begin by presenting the notation used in this paper. Vectors are denoted in bold, e.g., $\mathbf{x}\in\mathbb{R}^{n}$, and $\|\cdot\|$ denotes the Euclidean norm. For a differentiable function $J:\mathbb{R}^{n}\to\mathbb{R}$, its gradient and Hessian are denoted by $\nabla J(\mathbf{x})$ and $\nabla^{2}J(\mathbf{x})$, respectively. Inner products are written as $\langle\mathbf{u},\mathbf{v}\rangle=\mathbf{u}^{\top}\mathbf{v}$. For vectors (or blocks) $\mathbf{v}_{1},\dots,\mathbf{v}_{k}$, we use the stacking operator
$\operatorname{col}(\mathbf{v}_{1},\dots,\mathbf{v}_{k}):=\begin{bmatrix}\mathbf{v}_{1}^{\top}&\cdots&\mathbf{v}_{k}^{\top}\end{bmatrix}^{\top}$. (1)
We also make the following underlying assumptions that will be used throughout this work.
Assumption II.1.
The objective function $J(\mathbf{x})$ is continuously differentiable and has a locally Lipschitz continuous gradient.
Assumption II.2.
The optimization problem admits a finite optimal value, i.e., $J^{*}=\min_{\mathbf{x}}J(\mathbf{x})$.
These assumptions are used throughout this paper, even if not explicitly mentioned.
We emphasize that, under these standing assumptions, the convergence results established in this paper are generally local, unless stronger properties such as strong convexity or global convexity-concavity are imposed.
II-A Stationarity Vector
The central object in our framework is the stationarity vector $\mathbf{S}(\mathbf{z})\in\mathbb{R}^{n}$, which encodes the stationarity conditions of the optimization problem at hand. The form of $\mathbf{S}$ depends on the problem addressed (unconstrained, constrained, minimax, or GNE), where the vector $\mathbf{z}\in\mathbb{R}^{n}$ encodes the relevant variables which can be both the primal and dual variables and the dimension $n$ varies depending to the problem addressed. The unifying principle is that $\mathbf{S}(\mathbf{z}^{\star})=\mathbf{0}$ at any stationary point $\mathbf{z}^{\star}$.
Example II.3 (Stationarity in Unconstrained Optimization).
For the unconstrained minimization problem $\min_{\mathbf{x}}J(\mathbf{x})$, the stationarity vector reduces to the gradient, $\mathbf{S}(\mathbf{z})=\nabla J(\mathbf{x})$, where $\mathbf{z}=\mathbf{x}$. Hence, we consider the quadratic Lyapunov candidate
$V(\mathbf{x})=\tfrac{1}{2}\|\mathbf{S}(\mathbf{z})\|^{2}=\tfrac{1}{2}\|\nabla J(\mathbf{x})\|^{2},$ (2)
which vanishes precisely at stationary points.
Throughout the paper, feedback dynamics will be designed so that $\dot{V}(\mathbf{x})$ satisfies a predefined convergence law, such as exponential, FT, FxT, or PT decay, as detailed below.
II-B Desired Convergence Properties
We introduce a family of decay conditions for Lyapunov functions that generate different convergence behaviors. Rather than analyzing specific algorithms, we adopt a template-based viewpoint: given a candidate Lyapunov function $V$, the choice of a decay inequality $\dot{V}\leq-\sigma(V,t)$ determines the temporal properties of convergence. By varying the form of $\sigma$, one recovers exponential, FT, FxT, and PT regimes. These results are stated in the following lemmas for reference and will serve as fundamental ingredients throughout the remainder of the paper.
We summarize the known convergence results as lemmas to be used in later sections. These results are well established in control and optimization theory and are stated here for completeness.
Lemma II.4 (Exponential Stability [khalil_nonlinear_2002]).
If there exists $\alpha>0$ such that
$\dot{V}(\mathbf{z})\leq-\alpha V(\mathbf{z}),$ (3)
then $V(\mathbf{z})$ converges to zero exponentially, that is, $V(\mathbf{z}(t))\leq V(\mathbf{z}(0))e^{-\alpha t}$.
Lemma II.5 (FT Stability [bhat_finite-time_2000]).
If there exist $c>0$ and $\alpha\in(0,1)$ such that
$\dot{V}(\mathbf{z})\leq-cV(\mathbf{z})^{\alpha},$ (4)
then $V(\mathbf{z}(t))$ reaches zero in finite time, with settling-time estimate
$T(\mathbf{z}(0))\leq\frac{V(\mathbf{z}(0))^{1-\alpha}}{c(1-\alpha)}.$ (5)
Lemma II.6 (FxT Stability [polyakov_nonlinear_2012]).
If there exist $c_{1},c_{2}>0$, $\alpha\in(0,1)$, and $\beta>1$ such that
$\dot{V}(\mathbf{z})\leq-c_{1}V(\mathbf{z})^{\alpha}-c_{2}V(\mathbf{z})^{\beta},$ (6)
then $V(\mathbf{z}(t))$ converges to zero in fixed time, with a uniform upper bound
$T\leq\frac{1}{c_{1}(1-\alpha)}+\frac{1}{c_{2}(\beta-1)}.$ (7)
Lemma II.7 (PT Stability [song_time-varying_2017]).
If there exists a constant $c>0$ and a prescribed time $T>0$ such that
$\dot{V}(\mathbf{z})\;\leq\;-c\frac{T}{T-t}\,V(\mathbf{z}),\qquad t\in[0,T),$ (8)
Then $V(\mathbf{z}(t))$ converges to zero exactly at $t=T$, independent of the initial condition.
This work focuses on these four convergence regimes, but we note that other options exist for which the generalization of our framework to them is straightforward. For other convergence regimes, the reader is referred to [song_prescribed-time_2023].
The convergence laws introduced here are independent of any specific problem class. They serve as rate templates that can be combined with various Lyapunov constructions in the subsequent sections. In the following developments, these laws are employed to design optimizer dynamics for unconstrained and constrained optimization, minimax problems, and GNE seeking, where the desired convergence behavior is achieved by selecting the appropriate decay condition.
III Control-Centric Framework
This section develops a control-centric recipe for constructing continuous-time dynamics that solve optimization and equilibrium problems via Lyapunov methods. The key idea is to specialize a Lyapunov-driven design to a “plant” with ideal dynamics
$\dot{\mathbf{z}}\;=\;\mathbf{u},$ (9)
and to choose the Lyapunov function so that its set of minima coincides with the stationarity set of the optimization problem. For any candidate $V:\mathbb{R}^{n}\to\mathbb{R}_{\geq 0}$, the derivative along (9) is
$\dot{V}(\mathbf{z},t)\;=\;\nabla V(\mathbf{z})^{\top}\mathbf{u}(\mathbf{z},t).$ (10)
Our aim is to realize $\mathbf{u}(\mathbf{z},t)$ so that the decay inequality
$\dot{V}(\mathbf{z},t)\;\leq\;-\,\sigma\!\big(V(\mathbf{z}),t\big)$ (11)
holds, where the timing law $\sigma(\cdot,\cdot)$ encodes the desired convergence rate, e.g., exponential, FT, FxT, or PT, and will be instantiated via the rate lemmas collected in Sec. II-B.
III-A Optimization Lyapunov Function (OLF)
A central ingredient of our framework is the OLF which, analogously to Lyapunov functions for stability, certifies convergence of a dynamical optimizer to the stationarity set. This notion allows us to treat algorithm design as a control problem: the OLF encodes the desired optimality conditions, and the choice of decay law shapes the rate of convergence.
Definition III.1 (Optimization Lyapunov Function).
Let $\mathcal{S}\!=\!\{\mathbf{z}\in\mathbb{R}^{n}:\mathbf{S}(\mathbf{z})=\mathbf{0}\}$ denote the stationarity set. An optimization Lyapunov function is any continuously differentiable map $V:\mathbb{R}^{n}\!\to\!\mathbb{R}_{\geq 0}$ such that $V(\mathbf{z})=0$ if and only if $\mathbf{z}\in\mathcal{S}$.
III-B Nonsingularity Condition
Because $\mathbf{u}$ enters (10) through $\nabla V$, one cannot force $\dot{V}<0$ at states where $\nabla V=\mathbf{0}$ while $V>0$.
Assumption III.2 (Nonsingularity of The OLF Gradient).
Let $V(\mathbf{z})$ be an OLF as defined in III.1. Assume
$\nabla V(\mathbf{z})\;\neq\;\mathbf{0}$ (12)
whenever $V(\mathbf{z})\neq 0$.
If $\nabla V(\mathbf{z})=\mathbf{0}$ while $V(\mathbf{z})>0$, then (10) gives $\dot{V}(\mathbf{z},t)=0$ for all admissible $\mathbf{u}(\mathbf{z},t)$. Hence, no feedback based on $V$ can enforce $\dot{V}<0$ at that time. This assumption holds automatically in certain structured problems, such as a strongly convex cost function with convex inequality and affine equality constraints. We establish this formally in the ensuing results.
When the quadratic OLF, $V(\mathbf{z})=\tfrac{1}{2}\|\mathbf{S}(\mathbf{z})\|^{2}$, is used, the nonsingularity assumption becomes
$\nabla V(\mathbf{z})\;=\;\nabla\mathbf{S}(\mathbf{z})^{\top}\mathbf{S}(\mathbf{z})\;\neq\;\mathbf{0}$ (13)
whenever $\mathbf{S}(\mathbf{z})\neq\mathbf{0}$. Thus, the nonsingularity assumption becomes a nonorthogonality assumption as we assume in practice that
$\mathbf{S}(\mathbf{z})\;\notin\;\ker\big(\nabla\mathbf{S}(\mathbf{z})^{\top}\big).$ (14)
where $\ker(\cdot)$ denotes the null space. For the sake of generality and convenience of exposition, we will refer to this assumption as the nonsingularity assumption in the remainder of this paper.
III-C Feedback Designs Enforcing the Convergence Law
We present three constructive designs that enforce the decay law (11); they are ordered by restrictiveness on $\nabla V(\mathbf{z})$. Each can be paired with any admissible convergence law $\sigma(\cdot,\cdot)$.
Lemma III.3 (Hessian-Gradient Dynamics (HGD)).
Let $V(\mathbf{z})$ be an OLF. If Assumption III.2 holds along the trajectory, then the feedback
$\mathbf{u}(\mathbf{z},t)\;=\;-\,\frac{\sigma\!\big(V(\mathbf{z}),t\big)}{\|\nabla V(\mathbf{z})\|^{2}}\,\nabla V(\mathbf{z})$ (15)
enforces (11) with equality $\dot{V}=-\sigma(V,t)$. The convergence law satisfies
$\sigma(V,t)\;=\;\begin{cases}c\,V,&\text{(Exp.),}\\[4.0pt] k\,V^{\gamma},&\text{(FT), }\;0<\gamma<1,\\[4.0pt] a\,V^{\gamma}+b\,V^{\delta},&\text{(FxT), }\;0<\gamma<1<\delta,\\[4.0pt] \dfrac{\mu}{T-t}\,V,&\text{(PT), }\;t\in[0,T),\end{cases}$ (16)
with positive constants $a,b,c,k,\mu$ and $T$.
The term Hessian-gradient dynamics (HGD) is motivated by the unconstrained case: when $S(\mathbf{z})=\nabla J(\mathbf{x})$ with a quadratic Lyapunov function (2), the feedback involves the Hessian-gradient product $\nabla^{2}J(\mathbf{x})\,\nabla J(\mathbf{x})$. This yields a normalized second-order descent distinct from both gradient and Newton dynamics.
Proof.
Substitute (15) into (10) to get
$\dot{V}=\nabla V^{\top}\!\Big(-\,\tfrac{\sigma(V,t)}{\|\nabla V\|^{2}}\,\nabla V\Big)=-\,\sigma(V,t).$ (17)
The cases in Eq. (16) follow directly from Lemmas II.4-II.7. ∎
Example III.4 (HGD for Unconstrained Optimization).
Let $\mathbf{S}(\mathbf{x})=\nabla J(\mathbf{x})$ and $V(\mathbf{x})=\tfrac{1}{2}\|\mathbf{S}(\mathbf{x})\|^{2}$. Define the Hessian-gradient product
$\mathbf{w}(\mathbf{x})\;=\;\nabla^{2}J(\mathbf{x})\,\nabla J(\mathbf{x}).$ (18)
Assume $\mathbf{w}(\mathbf{x})\neq\mathbf{0}$ whenever $\nabla J(\mathbf{x})\neq\mathbf{0}$, i.e., that Assumption III.2 holds along the trajectory. Set the feedback
$\mathbf{u}(\mathbf{x},t)\;=\;-\,\frac{\sigma\!\big(V(\mathbf{x}),t\big)}{\|\mathbf{w}(\mathbf{x})\|^{2}}\,\mathbf{w}(\mathbf{x}),$ (19)
valid for any decay law $\sigma$; see Eq. (16) for the exponential, FT, FxT, and PT choices. Then along trajectories,
$\dot{V}\;=\;\nabla J^{\top}\nabla^{2}J\,\mathbf{u}\;=\;-\,\sigma\!\big(V(\mathbf{x}),t\big),$ (20)
so the timing law is enforced with equality.
If $J(\mathbf{x})=\tfrac{1}{2}\mathbf{x}^{\top}A\mathbf{x}-\mathbf{b}^{\top}\mathbf{x}$ with $A\succeq 0$ and $\mathbf{b}\notin\ker(A)$, then $\nabla J=A\mathbf{x}-\mathbf{b}$ and $\mathbf{w}=A\nabla J\neq\mathbf{0}$ whenever $\nabla J\neq\mathbf{0}$, so (19) yields (20) under any $\sigma$ in (16).
For clarity of exposition, the analysis adopts the standard quadratic Lyapunov candidate
$V(\mathbf{z})\;=\;\tfrac{1}{2}\big\|\mathbf{S}(\mathbf{z})\big\|^{2},$ (21)
which is used throughout the remainder of the paper.
Lemma III.5 (Newton Dynamics (ND)).
Suppose $\nabla\mathbf{S}(\mathbf{z})$ is invertible. Then
$\mathbf{u}(\mathbf{z},t)\;=\;-\,\lambda\!\big(V(\mathbf{z}),t\big)\,\nabla\mathbf{S}(\mathbf{z})^{-1}\,\mathbf{S}(\mathbf{z})$ (22)
satisfies (11) with equality when $\lambda(V,t)=\sigma(V,t)/(2V)$, where $\sigma(\cdot,\cdot)$ is chosen as presented in (16).
The term Newton Dynamics (ND) reflects that in the unconstrained case $S(\mathbf{z})=\nabla J(\mathbf{x})$ with a quadratic Lyapunov function (2), the feedback law reduces to the continuous-time Newton method $\dot{\mathbf{x}}=-\nabla^{2}J(\mathbf{x})^{-1}\nabla J(\mathbf{x})$ that is known to guarantee exponential convergence [brown_effective_1989].
Proof.
With $V=\tfrac{1}{2}\|\mathbf{S}\|^{2}$ and $\nabla V=\nabla\mathbf{S}^{\top}\mathbf{S}$,
$\dot{V}\;=\;-\,\lambda\,\mathbf{S}^{\top}\nabla\mathbf{S}\,\nabla\mathbf{S}^{-1}\mathbf{S}.$ (23)
Since $\nabla\mathbf{S}^{-1}$ is invertible, we obtain V˙= $\dot{V}=-\,2\lambda$ V. ∎
Example III.6 (ND for Unconstrained Optimization).
Let $\mathbf{S}(\mathbf{x})=\nabla J(\mathbf{x})$ and $V(\mathbf{x})=\tfrac{1}{2}\|\mathbf{S}(\mathbf{x})\|^{2}$. Assume $\nabla^{2}J(\mathbf{x})\succ 0$ for all $\mathbf{x}$, a sufficient condition for strict convexity. Set
$\mathbf{u}(\mathbf{x},t)\;=\;-\big[\nabla^{2}J(\mathbf{x})\big]^{-1}\,\nabla J(\mathbf{x})\,\lambda\!\big(V(\mathbf{x}),t\big),$ (24)
and choose $\lambda(V,t)=\sigma(V,t)/(2V)$, valid for any decay law $\sigma$, as shown in (16). Then
$\dot{V}=\nabla J^{\top}\nabla^{2}J\,\mathbf{u}=-\,\|\nabla J\|^{2}\,\lambda(V,t)=-\,\sigma\!\big(V(\mathbf{x}),t\big),$ (25)
so the timing law is enforced with equality.
If $J(\mathbf{x})=\tfrac{1}{2}\mathbf{x}^{\top}A\mathbf{x}-\mathbf{b}^{\top}\mathbf{x}$ with $A\succ 0$, then $\nabla J=A\mathbf{x}-\mathbf{b}$ and $\mathbf{u}=-A^{-1}\nabla J\,\lambda(V,t)$ satisfies (25) for any $\sigma$ in (16).
Lemma III.7 (Gradient Dynamics (GD)).
Suppose there exists $m>0$ such that for all $\mathbf{z}$,
$\frac{\nabla\mathbf{S}(\mathbf{z})+\nabla\mathbf{S}(\mathbf{z})^{\top}}{2}\succeq mI.$ (26)
Then the feedback
$\mathbf{u}(\mathbf{z},t)\;=\;-\,\gamma\!\big(V(\mathbf{z}),t\big)\,\mathbf{S}(\mathbf{z}),$ (27)
with $\gamma(V,t)=\sigma(V,t)/(2mV)$ satisfies
$\dot{V}\;\leq\;-\,\sigma(V,t),$ (28)
where $\sigma(\cdot,\cdot)$ is chosen as in (16).
The term Gradient Dynamics (GD) is chosen in analogy with the unconstrained case: when $S(\mathbf{z})=\nabla J(\mathbf{x})$ with a quadratic Lyapunov function (2), the feedback reduces to the classical gradient descent dynamics $\dot{\mathbf{x}}=-\nabla J(\mathbf{x})$ that is known to guarantee exponential convergence [brown_effective_1989].
Proof.
From (21) and (10),
$\dot{V}\;=\;-\,\gamma\,\mathbf{S}^{\top}\nabla\mathbf{S}\,\mathbf{S}.$ (29)
Decompose $\nabla\mathbf{S}$ as
$\nabla\mathbf{S}\;=\;H+K,\quad H:=\tfrac{1}{2}\!\big(\nabla\mathbf{S}+\nabla\mathbf{S}^{\top}\big),\;K:=\tfrac{1}{2}\!\big(\nabla\mathbf{S}-\nabla\mathbf{S}^{\top}\big).$
Here $K^{\top}=-K$, so $\mathbf{x}^{\top}K\mathbf{x}=0$ for any $\mathbf{x}$. Thus,
$\mathbf{S}^{\top}\nabla\mathbf{S}\,\mathbf{S}$ $=\mathbf{S}^{\top}H\mathbf{S}+\underbrace{\mathbf{S}^{\top}K\mathbf{S}}_{=\,0}=\mathbf{S}^{\top}H\mathbf{S}\;\geq\;m\|\mathbf{S}\|^{2}$ (30)
by (26). Substituting (30) into (29) yields
$\dot{V}$ $\;\leq\;-\,\gamma\,m\,\|\mathbf{S}\|^{2}=-\,\gamma\,m\,(2V)=-\,\sigma(V,t),$ (31)
which proves (28). ∎
Example III.8 (GD for Unconstrained Optimization).
Let $\mathbf{S}(\mathbf{x})=\nabla J(\mathbf{x})$ and define
$V(\mathbf{x})\;=\;\tfrac{1}{2}\|\mathbf{S}(\mathbf{x})\|^{2}.$ (32)
Assume $J$ is strongly convex: there exists $m>0$ such that
$\nabla^{2}J(\mathbf{x})\succeq mI\qquad\forall\,\mathbf{x}\in\mathbb{R}^{n}.$ (33)
Set the gradient feedback
$\mathbf{u}(\mathbf{x},t)\;=\;-\gamma\!\big(V(\mathbf{x}),t\big)\,\nabla J(\mathbf{x}),$ (34)
and choose $\gamma(V,t)=\sigma(V,t)/(2mV)$, valid for any decay law $\sigma$ as in (16). Then
$\dot{V}$ $=\nabla J^{\top}\nabla^{2}J\,\mathbf{u}\;=\;-\,\gamma\,\nabla J^{\top}\nabla^{2}J\,\nabla J$
$\leq-\,\gamma\,m\,\|\nabla J\|^{2}\;=\;-\,\sigma\!\big(V(\mathbf{x}),t\big),$ (35)
so the timing law is enforced in inequality form under strong convexity.
If $J(\mathbf{x})=\tfrac{1}{2}\mathbf{x}^{\top}A\mathbf{x}-\mathbf{b}^{\top}\mathbf{x}$ with $A\succeq mI$, then $\nabla J=A\mathbf{x}-\mathbf{b}$, and (34)-(35) hold for any $\sigma$ in (16).
Lemma III.3 is the least restrictive: it needs only $\nabla V\neq\mathbf{0}$ whenever $V\neq\mathbf{0}$ and it is valid for any OLF, whereas the other two are limited to the quadratic OLF (21). Lemma III.5 is stricter and implies nonsingularity (then $\mathbf{S}\in{\rm range}(\nabla\mathbf{S})$ so $\nabla V=\nabla\mathbf{S}^{\top}\mathbf{S}\neq\mathbf{0}$ for $\mathbf{S}\neq\mathbf{0}$). Lemma III.7 is strongest, requiring a uniform lower bound on the symmetric part $\tfrac{1}{2}(\nabla\mathbf{S}+\nabla\mathbf{S}^{\top})\succeq mI$ along the trajectory. The convergence law $\sigma(\cdot,\cdot)$ is left open here as a design choice as presented in (16).
Remark III.9 (Relation to Sontag’s Universal Construction).
Sontag’s classical work [sontag_universal_1989] considered nonlinear systems affine in the control and showed that, whenever a control-Lyapunov function exists, one can explicitly construct a feedback law that ensures asymptotic stabilization. The key idea is that the feedback recipe is universal: it applies to any system once a Lyapunov function is available. Our framework plays an analogous role in optimization: given an optimization Lyapunov function, the feedback designs in Lemmas III.3III.5, and III.7 provide universal constructions that not only ensure convergence but also shape its timing with exponential, FT, FxT, or PT guarantees.
III-D Numerical Example: Unconstrained Problem
We illustrate the three feedback laws on the $n$-dimensional, smooth, strongly convex model. We consider the classical log-sum-exp plus quadratic term function [castera_continuous_2024], for any $x\in\mathbb{R}^{n}$
$J(x)\;=\;\log\!\Bigg(\sum_{i=1}^{n}\big(e^{x_{i}}+e^{-x_{i}}\big)\Bigg)\;+\;\frac{1}{2}\,\|x\|^{2},$ (36)
which has the unique minimizer $x^{\star}=\mathbf{0}$. We compare the Hessian-gradient dynamics (HGD), the Newton dynamics (ND), and the gradient dynamics (GD). Each dynamics enforces the same Lyapunov decay template $\dot{V}=-\sigma(V,t)$, with $V(x)=\tfrac{1}{2}\|\nabla J(x)\|^{2}$. We test four convergence laws: exponential, FT, FxT, and PT, resulting in a total of twelve cases.
[…]
Figure 1 confirms that all $12$ cases realize the desired decay patterns induced by the law selector $\sigma(V,t)$. Table I reports the wall-clock costs, i.e., the actual CPU time required to run the simulations, whereas the times in Fig. 1 correspond to the continuous dynamics of (9) used in the simulation environment. For the exponential (Exp) case, the GD is fastest, with ND and HGD moderately slower. In the FT and FxT cases, HGD attains the lowest run times, while ND and GD are close behind. In the PT case, ND and HGD are more efficient than GD. Overall, the results show that while the three realizations (HGD/ND/GD) are all valid enforcers of the convergence laws, their runtime profiles differ depending on the chosen law: the HGD is usually the fastest except for Exp, where GD and ND are generally competitive.
Figure 1: Gradient norm trajectories $\|\nabla J(\mathbf{x})\|$ for $n=50$ under the exponential (Exp), finite-time (FT), fixed-time (FxT), and prescribed-time (PT) laws, denoted by blue, red, magenta, and black lines, respectively. Three realizations are compared: GD, ND, and HGD, denoted by solid, dashed, and dashed-dotted lines, respectively. The test function is the log-sum-exp plus quadratic model (36), initialized at $\mathbf{x}_{0}=[1,\ldots,1]^{\top}$.
TABLE I: Wall-clock CPU times (ms) for the $n=50$ log-sum-exp example under the four convergence laws and three feedback realizations.
Case Exp FT FxT PT
ND 132 58 79 92
GD 73 48 77 134
HGD 86 19 51 81
All three feedback laws achieve convergence with the selected timing law, but differ in their computational profiles. The HGD design avoids matrix inversion at the cost of additional multiplications; the ND is fastest under strict convexity, and the GD applies whenever strong convexity holds. These examples demonstrate how the OLF framework accommodates different structural assumptions while unifying exponential, FT, FxT, and PT guarantees. In the next section, we extend the approach to constrained optimization problems, where the stationarity vector incorporates equality and inequality constraints.
We note that the OLF provides a template for unifying algorithm design across problem classes. Given an optimization problem, one first identifies the appropriate stationarity conditions and encodes them in $V$. Next, by selecting a convergence law, a feedback law can be constructed that enforces the desired rate of decay. In the following sections, we apply this methodology to constrained optimization, minimax formulations, and GNE seeking problems.
IV Constrained Optimization
We now extend the control-centric framework to constrained optimization problems that include equality and inequality constraints, covering exponential, FT, FxT, and PT convergence laws. For simplicity of exposition, we restrict ourselves to the Hessian-gradient dynamics (HGD) case, as it requires the least restrictive assumptions. Note that generalizing the derivations of this section and the following two for the ND and GD cases is straightforward.
IV-A Problem Setup and Preliminaries
Consider the constrained problem
$\min_{\mathbf{x}\in\mathbb{R}^{n}}$~ $J(\mathbf{x})$ (37a)
s.t. $\mathbf{g}(\mathbf{x})\leq\mathbf{0},$ (37b)
$\mathbf{h}(\mathbf{x})=\mathbf{0},$ (37c)
with the associated Lagrangian
$L(\mathbf{x},\boldsymbol{\lambda},\boldsymbol{\mu})=J(\mathbf{x})+\boldsymbol{\lambda}^{\top}\mathbf{g}(\mathbf{x})+\boldsymbol{\mu}^{\top}\mathbf{h}(\mathbf{x}),$ (38)
where $J:\mathbb{R}^{n}\to\mathbb{R}$, $\mathbf{g}:\mathbb{R}^{n}\to\mathbb{R}^{p}$, $\mathbf{h}:\mathbb{R}^{n}\to\mathbb{R}^{q}$, $\boldsymbol{\lambda}\in\mathbb{R}^{p}_{\geq 0}$, and $\boldsymbol{\mu}\in\mathbb{R}^{q}$.
Definition IV.1 (KKT conditions).
A point $(\mathbf{x}^{\star},\boldsymbol{\lambda}^{\star},\boldsymbol{\mu}^{\star})$ satisfies the KKT conditions if
$\nabla J(\mathbf{x}^{\star})+\nabla\mathbf{g}(\mathbf{x}^{\star})^{\top}\boldsymbol{\lambda}^{\star}+\nabla\mathbf{h}(\mathbf{x}^{\star})^{\top}\boldsymbol{\mu}^{\star}=\mathbf{0},$ (39a)
$\mathbf{g}(\mathbf{x}^{\star})\leq\mathbf{0},~~\boldsymbol{\lambda}^{\star}\geq\mathbf{0},~~\langle\boldsymbol{\lambda}^{\star},\mathbf{g}(\mathbf{x}^{\star})\rangle=0,$ (39b)
$\mathbf{h}(\mathbf{x}^{\star})=\mathbf{0}.$ (39c)
IV-B Fischer-Burmeister Function and Its Usage for Stationarity
We encode KKT conditions associated with the inequality constraints in (39b) via the Fischer-Burmeister (FB) function [fischer_special_1992]:
$\phi(a,b):=\sqrt{a^{2}+b^{2}}-(a+b).$ (40)
Then $\boldsymbol{\phi}(\boldsymbol{\lambda},\mathbf{g})=0$ if and only if $\boldsymbol{\lambda}\geq\mathbf{0}$, $\mathbf{g}\leq\mathbf{0}$, and $\boldsymbol{\lambda}^{\top}\mathbf{g}=0$, where $\phi$ applies elementwise;
The FB function is nonsmooth near the origin, so we use its smoothed version [liao-mcpherson_regularized_2019]
$\phi_{\varepsilon}(a,b):=\sqrt{a^{2}+b^{2}+\varepsilon^{2}}-(a+b),\quad\varepsilon>0,$ (41)
where the smoothing satisfies
$\left\|\boldsymbol{\phi}_{\varepsilon}(\boldsymbol{\lambda},\mathbf{g})-\boldsymbol{\phi}(\boldsymbol{\lambda},\mathbf{g})\right\|\leq\varepsilon,$ (42)
for any $\varepsilon>0$. The smoothed FB mapping ensures the feasibility of the inequality constraints up to $\varepsilon>0$, yielding an $\varepsilon$-KKT solution. One may select $\varepsilon$ arbitrarily small to approach exact feasibility while preserving differentiability and good numerical properties [liao-mcpherson_regularized_2019]. If exact KKT is required, one may instead use the stationarity vector presented in Remark IV.2 below.
Define the stacked variable $\mathbf{z}:=\mathrm{col}(\mathbf{x},\boldsymbol{\lambda},\boldsymbol{\mu})\in\mathbb{R}^{n+p+q}$ and the stationarity vector
$\mathbf{S}(\mathbf{z})=\begin{bmatrix}\nabla_{\mathbf{x}}L(\mathbf{x},\boldsymbol{\lambda},\boldsymbol{\mu})\\[2.0pt] \boldsymbol{\phi}_{\varepsilon}\!\big(\boldsymbol{\lambda},\mathbf{g}(\mathbf{x})\big)\\[2.0pt] \mathbf{h}(\mathbf{x})\end{bmatrix}$. (43)
Then $\mathbf{S}(\mathbf{z})=\mathbf{0}$ enforces (39a)-(39c) up to the smoothing tolerance $\varepsilon$.
Remark IV.2 (Alternative Stationarity Without Smoothing).
If exact feasibility is required, one may use the alternative stationarity vector
$\mathbf{S}(\mathbf{z})=\begin{bmatrix}\nabla_{\mathbf{x}}L(\mathbf{x},\boldsymbol{\lambda},\boldsymbol{\mu})\\ \boldsymbol{\lambda}^{\top}\mathbf{g}(\mathbf{x})\\ \mathbf{g}_{+}(\mathbf{x})\\ \boldsymbol{\lambda}_{-}\\ \mathbf{h}(\mathbf{x})\end{bmatrix}$, (44)
with $\mathbf{g}_{+}(\mathbf{x})=\max(\mathbf{g}(\mathbf{x}),\mathbf{0})$ and $\boldsymbol{\lambda}_{-}=\max(-\boldsymbol{\lambda},\mathbf{0})$ applied componentwise. In this case, $\mathbf{S}(\mathbf{z})$ involves max operators and is therefore nonsmooth. However, the quadratic OLF $V(\mathbf{z})=\tfrac{1}{2}\|\mathbf{S}(\mathbf{z})\|^{2}$ remains continuously differentiable [rockafellar_dual_1973].
We also include the following classical regularity assumptions for constrained optimization [beck_introduction_2014].
Assumption IV.3 (LICQ).
Let $\mathcal{A}(\mathbf{x}):=\{\,i\in\{1,\ldots,p\}\mid g_{i}(\mathbf{x})=0\,\}$ denote the active set. We say the linear independence constraint qualification (LICQ) holds at $\mathbf{x}$ if the set
$\Big\{\,\nabla h_{j}(\mathbf{x})~:~j=1,\ldots,q\,\Big\}\cup\Big\{\,\nabla g_{i}(\mathbf{x})~:~i\in\mathcal{A}(\mathbf{x})\,\Big\}$ (45)
is linearly independent.
Assumption IV.4 (Generalized Slater).
For the inequality constraints $\mathbf{g}(\mathbf{x})\leq\mathbf{0}$, generalized Slater’s condition holds if there exists $\bar{\mathbf{x}}$ such that $\mathbf{h}(\bar{\mathbf{x}})=\mathbf{0}$ and $\mathbf{g}(\bar{\mathbf{x}})<\mathbf{0}$.
IV-C Results
We implement the Hessian-gradient dynamics (HGD):
$\dot{\mathbf{z}}=\mathbf{u}(\mathbf{z},t)=-\frac{\nabla\mathbf{S}(\mathbf{z})^{\top}\mathbf{S}(\mathbf{z})}{\left\|\nabla\mathbf{S}(\mathbf{z})^{\top}\mathbf{S}(\mathbf{z})\right\|^{2}}\;\sigma\!\big(V(\mathbf{z}),t\big),$ (46)
where $\sigma(\cdot,\cdot)$ is the scalar law selector. This feedback yields
$\dot{V}=-\,\sigma\!\big(V(\mathbf{z}),t\big),$ (47)
on the domain where $\nabla\mathbf{S}(\mathbf{z})^{\top}\mathbf{S}(\mathbf{z})\neq\mathbf{0}$.
We now present the theorem for the constrained case, which covers both equality and inequality constraints using the HGD law. It recovers the desired convergence by using (16).
Theorem IV.5 (HGD for Constrained Optimization).
Consider (37) with $J,\mathbf{g},\mathbf{h}\in\mathcal{C}^{2}$, and define $\mathbf{S}$ via (43). Fix $\varepsilon>0$ and assume that Assumption III.2 holds throughout the entire trajectory of $\mathbf{z}(t)$ and Assumption IV.3 holds at the KKT point. Then, using the HGD dynamics (46), with the convergence profile implied by $\sigma(\cdot,\cdot)$ as presented in (16), the KKT conditions are enforced up to the smoothing tolerance $\varepsilon$.
Proof.
Assumption IV.3 guarantees the existence of the multipliers [beck_introduction_2014, Thm. 11.6]. By Assumption III.2, $\nabla\mathbf{S}(\mathbf{z})^{\top}\mathbf{S}(\mathbf{z})\neq\mathbf{0}$ whenever $\mathbf{S}(\mathbf{z})\neq\mathbf{0}$, so the denominator in (46) is well-defined. Using Eq. (46), we get
$\dot{V}=\mathbf{S}^{\top}\nabla\mathbf{S}\left(-\frac{\nabla\mathbf{S}^{\top}\mathbf{S}}{\|\nabla\mathbf{S}^{\top}\mathbf{S}\|^{2}}\sigma(V,t)\right)=-\,\sigma(V,t),$ (48)
which is (47). Since $V=\frac{1}{2}\|\mathbf{S}\|^{2}$, the decay of $V$ to zero implies $\mathbf{S}(\mathbf{z}(t))\to\mathbf{0}$ with the indicated timing; feasibility of the inequality constraints is enforced up to the smoothing tolerance $\varepsilon$ of (41). ∎
Next we demonstrate that uniqueness and nonsingularity hold under strong convexity, affine equalities, and convex inequalities.
Corollary IV.6 (Global Optimality).
Consider (37) with $J,\mathbf{g},\mathbf{h}\in\mathcal{C}^{2}$, and define $\mathbf{S}$ via (43). Fix $\varepsilon>0$ and assume that there exists $\bar{\mathbf{x}}$ as shown in Assumption IV.4 and that Assumption IV.3 holds at the KKT point. Suppose $J$ is strongly convex, the equality constraints are affine $\mathbf{h}(\mathbf{x})=\mathbf{A}\mathbf{x}-\mathbf{b}$ with $\mathbf{A}$ full row rank, and each $g_{i}$ is convex. Then, the KKT point $(\mathbf{x}^{\star},\boldsymbol{\lambda}^{\star},\boldsymbol{\mu}^{\star})$ is unique and is the unique global minimizer. Hence, Theorem IV.5 ensures global convergence (with the chosen timing law) to the unique solution.
Strong convexity and LICQ imply uniqueness of the KKT solution, while Slater’s condition ensures feasibility. Moreover, there is no need to assume nonsingularity as it is guaranteed in this case. Together, these rule out spurious stationary points of $V(\mathbf{z})$. A complete argument is provided in Appendix A-A.
Remark IV.7 (On Radial Unboundedness).
Unlike standard Lyapunov-based stability analyses, the result in Corollary IV.6 does not require the Lyapunov function to be radially unbounded in $\mathbf{z}$. This is because the dynamics are expressed in feedback form (46), which directly enforces boundedness and drives the trajectories toward the unique KKT point. Under the strong convexity and convex feasibility assumptions, this point coincides with the global optimizer. Hence, global convergence follows from the closed-loop construction itself, rather than from the growth of $V$ at infinity.
IV-D Illustrative Example: Network Utility Maximization
Consider the classical network utility maximization (NUM) problem, widely used in the study of continuous-time primal-dual dynamics [feijer_stability_2010]. Let $\mathbf{x}\in\mathbb{R}_{\geq 0}^{S}$ denote the vector of source rates, $\mathbf{R}\in\{0,1\}^{L\times S}$ the routing matrix (with $R_{ij}=1$ if source $j$ uses link $i$), and $\mathbf{c}\in\mathbb{R}^{L}_{>0}$ the vector of link capacities. Each source $j$ has a strictly concave, continuously differentiable utility $U_{j}:\mathbb{R}_{>0}\to\mathbb{R}$ associated with it. The NUM problem reads
$\max_{\mathbf{x}\,>\,\mathbf{0}}\;\;\sum_{j=1}^{S}U_{j}(x_{j})\quad\text{s.t.}\quad\mathbf{R}\mathbf{x}\,\leq\,\mathbf{c}.$ (49)
Under standard assumptions (strict concavity of $U_{j}$ and Slater’s condition), (49) has a unique primal optimum. We set $U_{j}=\alpha_{j}\log(x_{j})$, a strictly concave function for $x_{j}>0$, where $\alpha_{j}>0$.
Problem (49) fits directly into our constrained setup. In particular, the quadratic OLF (21) satisfies the decay condition (47) under each of the decay laws in (16). We use the stationarity vector from (43) with $\varepsilon=10^{-6}$ for the smoothed FB.
Classical primal-dual dynamics for (49) exhibit asymptotic or exponential convergence under strict concavity [feijer_stability_2010]. Within our framework, the same problem admits FT, FxT, and PT guarantees, tightening classical results while preserving the problem structure.
Figure 2 confirms that the four timing laws realize the expected decay patterns: exponential trajectories converge asymptotically, while the FT and FxT designs reach zero in bounded time. The PT law enforces convergence exactly at the user-specified horizon. Table II highlights the corresponding computational cost. All four laws terminate within comparable runtimes, with the PT case being the most efficient in this experiment. These results demonstrate that our law selector carries over unchanged to a canonical constrained problem, providing FT, FxT, and PT guarantees beyond the classical exponential behavior reported in [feijer_stability_2010].
Figure 2: Decay of the optimality Lyapunov function $V(z(t))=\tfrac{1}{2}\|S(z(t))\|^{2}$ for the Network Utility Maximization (NUM) problem under the four convergence laws. Exp converges asymptotically, FT and FxT converge in bounded time, and PT enforces convergence exactly at the prescribed horizon.
TABLE II: Wall-clock CPU time (ms) for the NUM problem.
Exp FT FxT PT
239 218 324 168
Remark IV.8 (Stationarity monitoring).
The stationarity vector $\mathbf{S}(\mathbf{z})$ in (43) stacks the Lagrangian gradient and the smoothed FB, i.e.,
$\mathbf{S}(\mathbf{z})=\begin{bmatrix}\nabla_{\mathbf{x}}L(\mathbf{x},\boldsymbol{\lambda})\\[2.0pt] \boldsymbol{\phi}_{\varepsilon}(\boldsymbol{\lambda},\mathbf{R}\mathbf{x}-\mathbf{c})\end{bmatrix}$. (50)
For any vector norm, the norm of the stacked vector dominates the norm of each block. Therefore,
$\|\mathbf{S}(\mathbf{z})\|\;\geq\;\max\!\Big\{\,\|\nabla_{\mathbf{x}}L(\mathbf{x},\boldsymbol{\lambda})\|,\;\|\boldsymbol{\phi}_{\varepsilon}(\boldsymbol{\lambda},\mathbf{R}\mathbf{x}-\mathbf{c})\|\,\Big\}.$ (51)
Thus, driving $V(\mathbf{z})=\tfrac{1}{2}\|\mathbf{S}(\mathbf{z})\|^{2}$ below a threshold automatically guarantees that each optimality component is satisfied to the same numerical accuracy. This allows a single scalar Lyapunov function to monitor convergence of all optimality conditions.
This section showed how equality and inequality constraints can be incorporated via a unified stationarity mapping and the HGD feedback law. The framework recovers exponential, FT, FxT, and PT convergence guarantees in constrained problems. In the next section, we extend this approach to constrained minimax formulations, where two agents interact under shared feasibility constraints.
V Constrained Minimax Problems
Constrained minimax problems are central in robust control, adversarial machine learning, and game-theoretic models of resource allocation, where two agents with conflicting objectives interact under shared feasibility constraints. Classical minimax results [kose_solutions_1956, cherukuri_saddle-point_2017] establish asymptotic or exponential convergence for convex-concave problems with constraints. More recent works demonstrate FT and FxT convergence in unconstrained settings [romero_finite-time_2020, garg_fixed-time_2021, ozaslan_exponential_2024]. However, convergence in FT, FxT, and PT for constrained convex-concave problems has not been documented. This section addresses this gap using our control-centric framework with the HGD case.
V-A Problem Statement and Existence
We consider the constrained minimax problem
$\min_{\mathbf{x}\in\mathbb{R}^{n_{x}}}\ \max_{\mathbf{y}\in\mathbb{R}^{n_{y}}}\quad$ $J(\mathbf{x},\mathbf{y})$ (52a)
s.t. $A\,\mathbf{x}+B\,\mathbf{y}-\mathbf{b}=\mathbf{0},$ (52b)
$\mathbf{G}(\mathbf{x},\mathbf{y})\leq\mathbf{0},$ (52c)
where $J(\cdot,\mathbf{y})$ is convex in $\mathbf{x}$ for each fixed $\mathbf{y}$, and $J(\mathbf{x},\cdot)$ is concave in $\mathbf{y}$ for each fixed $\mathbf{x}$. The matrices $A\in\mathbb{R}^{q\times n_{x}}$ and $B\in\mathbb{R}^{q\times n_{y}}$ define $q$ affine equality constraints, and $\mathbf{G}:\mathbb{R}^{n_{x}+n_{y}}\to\mathbb{R}^{m}$ collects $m$ convex inequality constraints.
Standard regularity assumptions, such as LICQ in IV.3 and Slater in IV.4, are assumed. By Sion’s minimax theorem [sion_general_1958], a saddle point $(\mathbf{x}^{\star},\mathbf{y}^{\star})$ exists and $\min\max=\max\min$ holds for the convex-concave case. We also use the nonsingularity requirement already stated in Assumption III.2 of the manuscript.
The KKT conditions for (52) introduce multipliers $\boldsymbol{\mu}\in\mathbb{R}^{q}$ for the equalities (52b) and $\boldsymbol{\lambda}\in\mathbb{R}^{m}_{\geq 0}$ for the inequalities (52c). The KKT system reads
$\nabla_{\mathbf{x}}J(\mathbf{x}^{\star},\mathbf{y}^{\star})+A^{\top}\boldsymbol{\mu}^{\star}+\nabla_{\mathbf{x}}\mathbf{G}(\mathbf{x}^{\star},\mathbf{y}^{\star})^{\top}\boldsymbol{\lambda}^{\star}=0,$ (53a)
- $\nabla_{\mathbf{y}}J(\mathbf{x}^{\star},\mathbf{y}^{\star})+B^{\top}\boldsymbol{\mu}^{\star}+\nabla_{\mathbf{y}}\mathbf{G}(\mathbf{x}^{\star},\mathbf{y}^{\star})^{\top}\boldsymbol{\lambda}^{\star}=0,$ (53b)
$\mathbf{G}(\mathbf{x}^{\star},\mathbf{y}^{\star})\leq 0,\quad\boldsymbol{\lambda}^{\star}\geq 0,\quad\langle\boldsymbol{\lambda}^{\star},\mathbf{G}(\mathbf{x}^{\star},\mathbf{y}^{\star})\rangle=0,$ (53c)
$A\mathbf{x}^{\star}+B\mathbf{y}^{\star}-\mathbf{b}=0.$ (53d)
Let $\mathbf{z}\coloneqq(\mathbf{x},\mathbf{y},\boldsymbol{\lambda},\boldsymbol{\mu})$, where $\boldsymbol{\lambda}$ and $\boldsymbol{\mu}$ are the Lagrange multipliers for the inequality and equality constraints in (52). We define the stationarity vector $\mathbf{S}(\mathbf{z})$ as
$S(\mathbf{z})=\begin{bmatrix}\nabla_{\mathbf{x}}J(\mathbf{x},\mathbf{y})+A^{\top}\boldsymbol{\mu}+\nabla_{\mathbf{x}}\mathbf{G}(\mathbf{x},\mathbf{y})^{\top}\boldsymbol{\lambda}\\[2.0pt] -\nabla_{\mathbf{y}}J(\mathbf{x},\mathbf{y})+B^{\top}\boldsymbol{\mu}+\nabla_{\mathbf{y}}\mathbf{G}(\mathbf{x},\mathbf{y})^{\top}\boldsymbol{\lambda}\\[2.0pt] \boldsymbol{\phi}_{\varepsilon}\!\big(\boldsymbol{\lambda},\,\mathbf{G}(\mathbf{x},\mathbf{y})\big)\\[2.0pt] A\mathbf{x}+B\mathbf{y}-\mathbf{b}\end{bmatrix}$. (54)
The condition $\mathbf{S}(\mathbf{z})=\mathbf{0}$ is equivalent to the KKT system in (53), encompassing stationarity in both $\mathbf{x}$ and $\mathbf{y}$ together with feasibility of the constraints up to $\varepsilon>0$.
Following our framework, we associate to $\mathbf{S}$ the quadratic OLF (21), and define the dynamics
$\dot{\mathbf{z}}\;=\;\mathbf{u}(\mathbf{z},t)\;=\;-\,\frac{\nabla\mathbf{S}(\mathbf{z})^{\top}\mathbf{S}(\mathbf{z})}{\big\|\nabla\mathbf{S}(\mathbf{z})^{\top}\mathbf{S}(\mathbf{z})\big\|^{2}}\;\sigma\!\big(V(\mathbf{z}),t\big).$ (55)
This ensures that the derivative of $V$ along trajectories is shaped by the selected decay law.
V-B Convex-Concave Case
Theorem V.1 (Convex-Concave Minimax).
Suppose $J(\mathbf{x},\mathbf{y})\in\mathcal{C}^{2}$ is convex in $\mathbf{x}$ and concave in $\mathbf{y}$, and the feasible set of (52) is compact. Assume that Assumption III.2 holds throughout the entire trajectory of $\mathbf{z}(t)$, Assumption IV.3 holds at the KKT point, and that there exists $(\bar{\mathbf{x}},\bar{\mathbf{y}})$ as shown in Assumption IV.4. Then, a saddle point $(\mathbf{x}^{\star},\mathbf{y}^{\star})$ exists, and the HGD dynamics in (55) converges to the saddle point according to the chosen convergence rate, as presented in (16).
Proof.
By Sion’s minimax theorem [sion_general_1958], a saddle point $(\mathbf{x}^{\star},\mathbf{y}^{\star})$ exists. Along the dynamics (55), the derivative of $V$ satisfies
$\dot{V}(\mathbf{z}(t))=-\,\sigma\!\big(V(\mathbf{z}(t)),t\big),$ (56)
so $V(\mathbf{z}(t))\to 0$ with the chosen timing law, which implies $S(\mathbf{z}(t))\to\mathbf{0}$. ∎
Classical minimax dynamics have established asymptotic or exponential convergence under convex-concave assumptions [feijer_stability_2010, cherukuri_saddle-point_2017]. Our framework recovers these results as special cases and further guarantees FT, FxT, or PT convergence by an appropriate choice of the decay law.
V-C Strongly Convex-Strongly Concave Case
We next consider the standard strengthening that yields uniqueness.
Corollary V.2 (Strongly Convex-Strongly Concave).
Suppose $J(\mathbf{x},\mathbf{y})\in\mathcal{C}^{2}$ is $m_{x}$-strongly convex in $\mathbf{x}$ and $J(\mathbf{x},\cdot)$ is $m_{y}$-strongly concave in $\mathbf{y}$ with $m_{x},m_{y}>0$. Assume Assumption IV.3 holds at the KKT point and that there exists $(\bar{\mathbf{x}},\bar{\mathbf{y}})$ as shown in Assumption IV.4. Then the saddle point $(\mathbf{x}^{\star},\mathbf{y}^{\star})$ of (52) is unique, and the trajectories of (55) converge to $(\mathbf{x}^{\star},\mathbf{y}^{\star})$ with the selected convergence guarantee, as presented in (16).
Similarly to the constrained optimization case, when strong convexity/concavity holds, we get automatic nonsingularity of Assumption III.2. Moreover, as discussed in Remark IV.7, radial unboundedness of the Lyapunov function is not required, since the saddle-point dynamics are expressed in feedback form and inherently ensure bounded trajectories through the closed-loop structure.
Proof.
By Sion’s theorem, a saddle point exists for (52). Strong convexity in $\mathbf{x}$ and strong concavity in $\mathbf{y}$ exclude multiple saddle points, hence uniqueness. These properties imply Assumption III.2 holds automatically in this setting. The Lyapunov analysis with quadratic $V$ and dynamics (55) then yields convergence to $(\mathbf{x}^{\star},\mathbf{y}^{\star})$ with the chosen timing law. ∎
Classical constrained minimax dynamics achieve exponential convergence for convex-concave problems [feijer_stability_2010, cherukuri_saddle-point_2017]. Recent FT/FxT solutions address unconstrained settings [romero_finite-time_2020, garg_fixed-time_2021, ozaslan_exponential_2024]. Theorem V.1 and Corollary V.2 extend FT/FxT/PT guarantees to constrained convex-concave minimax problems.
This section establishes convergence for constrained convex-concave minimax problems according to a chosen convergence rate, extending beyond existing asymptotic or exponential guarantees in the literature. These results illustrate how the OLF framework generalizes across problem classes with shared feasibility constraints. In the next section, we address the GNE seeking problems, where multiple agents interact strategically under such shared constraints.
VI Generalized Nash Equilibrium Seeking Problems
Generalized Nash equilibrium (GNE) seeking problems arise when agents share coupling constraints, such as common resources or safety requirements, while pursuing individual objectives. These problems have been extensively studied from both the variational inequality and optimization perspectives; see the comprehensive survey in [facchinei_generalized_2010]. Classical reformulations based on Nikaido-Isoda functions and their regularized variants [von_heusinger_optimization_2009] provide optimization-based characterizations of (normalized) GNE. Algorithmic approaches have primarily relied on penalty and augmented Lagrangian methods, which embed the constraints into the agents’ costs. Representative contributions include penalty reformulations [facchinei_penalty_2010], augmented Lagrangian frameworks [kanzow_augmented_2016], and, more recently, continuous-time penalty dynamics that establish asymptotic or exponential convergence [sun_continuous-time_2021].
Despite these advances, existing continuous-time GNE seeking algorithms remain fundamentally penalty- or projection-based and do not provide guarantees of FT, FxT, or PT convergence. Moreover, their convergence analyses hinge on asymptotic Lyapunov arguments. In contrast, the control-centric OLF framework developed in this paper provides explicit convergence rate guarantees that extend beyond asymptotic or exponential stability. In what follows, we specialize this framework to GNE problems with strongly monotone games with convex shared constraints.
VI-A Problem statement and stationarity vector
Consider $N$ players, where each player $i\in\{1,\dots,N\}$ chooses a decision vector $\mathbf{x}_{i}\in\mathbb{R}^{n_{i}}$ subject to its local constraint set $\mathcal{X}_{i}$. Let $\mathbf{x}_{-i}$ denote the stacked decisions of all other players. Then, each player seeks to minimize its local cost function
$\min_{\mathbf{x}_{i}\in\mathcal{X}_{i}}J_{i}(\mathbf{x}_{i},\mathbf{x}_{-i}),$ (57)
where $J_{i}:\mathbb{R}^{n}\!\to\!\mathbb{R}$ is continuously differentiable and depends on both the player’s own decision and those of the others.
A profile $\bar{\mathbf{x}}=\operatorname{col}(\bar{\mathbf{x}}_{1},\dots,\bar{\mathbf{x}}_{N})$ with $n=\sum_{i}n_{i}$ is a GNE if it is feasible for the shared constraints and no player can reduce its cost by a unilateral deviation that also respects those shared constraints; namely,
$J_{i}(\bar{\mathbf{x}}_{i},\bar{\mathbf{x}}_{-i})\leq J_{i}(\mathbf{x}_{i},\bar{\mathbf{x}}_{-i})\quad\forall\,\mathbf{x}_{i}\in\mathcal{X}_{i}(\bar{\mathbf{x}}_{-i}),$ (58)
with $\mathcal{X}_{i}(\mathbf{x}_{-i}):=\{\mathbf{x}_{i}\mid\mathbf{A}\mathbf{x}=\mathbf{b},\ \mathbf{g}(\mathbf{x})\leq\mathbf{0}\}$ where $\mathbf{A}\in\mathbb{R}^{q\times n}$, $\mathbf{b}\in\mathbb{R}^{q}$, and $\mathbf{g}:\mathbb{R}^{n}\to\mathbb{R}^{m}$ is $C^{1}$. This differs from a standard Nash equilibrium because each player’s feasible set depends on the others via the shared constraints. In convex settings, equilibria for which all players share the same multipliers $(\boldsymbol{\lambda},\boldsymbol{\mu})$ for the joint constraints are called variational GNE (v-GNE); these coincide with solutions of a single KKT system [dreves_solution_2011] and are the natural target for our OLF design.
Denote the (stacked) pseudogradient mapping
$\mathcal{G}(\mathbf{x})\coloneqq\operatorname{col}\!\big(\nabla_{\mathbf{x}_{1}}J_{1}(\mathbf{x}_{1},\mathbf{x}_{-1}),\dots,\nabla_{\mathbf{x}_{N}}J_{N}(\mathbf{x}_{N},\mathbf{x}_{-N})\big).$ (59)
We adopt the variational formulation of GNE, using shared Lagrange multipliers $\boldsymbol{\lambda}\in\mathbb{R}^{m}_{\geq 0}$ for $\mathbf{g}(\mathbf{x})\leq\mathbf{0}$ and $\boldsymbol{\mu}\in\mathbb{R}^{q}$ for $\mathbf{A}\mathbf{x}=\mathbf{b}$. The KKT system reads
$\mathcal{G}(\mathbf{x})+\nabla\mathbf{g}(\mathbf{x})^{\top}\boldsymbol{\lambda}+\mathbf{A}^{\top}\boldsymbol{\mu}=\mathbf{0},\quad\mathbf{A}\mathbf{x}-\mathbf{b}=\mathbf{0},$ (60a)
$\mathbf{g}(\mathbf{x})\leq\mathbf{0},\ \boldsymbol{\lambda}\geq\mathbf{0},\ \langle\boldsymbol{\lambda},\mathbf{g}(\mathbf{x})\rangle=0.$ (60b)
Following our convention of encoding inequality constraints via the smoothed FB function, we define
$\mathbf{r}_{\mathrm{eq}}(\mathbf{x})\coloneqq\mathbf{A}\mathbf{x}-\mathbf{b},$ (61)
$\mathbf{r}_{\mathrm{ineq}}(\mathbf{x},\boldsymbol{\lambda})\coloneqq\boldsymbol{\phi}_{\varepsilon}\!\big(\boldsymbol{\lambda},\,\mathbf{g}(\mathbf{x})\big).$ (62)
With the stacked primal-dual variable $\mathbf{z}\coloneqq\operatorname{col}(\mathbf{x},\boldsymbol{\lambda},\boldsymbol{\mu})$, the stationarity vector is
$\mathbf{S}(\mathbf{z})\coloneqq\operatorname{col}\!\Big(\mathcal{G}(\mathbf{x})+$ $\nabla\mathbf{g}(\mathbf{x})^{\top}\boldsymbol{\lambda}+\mathbf{A}^{\top}\boldsymbol{\mu},\;$
$\mathbf{r}_{\mathrm{eq}}(\mathbf{x}),\;\mathbf{r}_{\mathrm{ineq}}(\mathbf{x},\boldsymbol{\lambda})\Big).$ (63)
Then $\mathbf{S}(\mathbf{z}^{\star})=\mathbf{0}$ if and only if $\mathbf{z}^{\star}$ satisfies (60a) exactly and (60b) up to $\varepsilon>0$.
We define the following underlying assumption for the GNE problem.
Assumption VI.1 (Strong monotonicity).
There exists $m>0$ such that for all $\mathbf{x},\mathbf{y}\in\mathbb{R}^{n}$
$\langle\mathcal{G}(\mathbf{x})-\mathcal{G}(\mathbf{y}),\,\mathbf{x}-\mathbf{y}\rangle\geq m\|\mathbf{x}-\mathbf{y}\|^{2}.$ (64)
Moreover, $\mathcal{G}$ is locally Lipschitz, $\mathbf{A}$ has full row rank, $\mathbf{g}$ is $C^{1}$, and the shared feasible set $\{\mathbf{x}:\mathbf{A}\mathbf{x}=\mathbf{b},\,\mathbf{g}(\mathbf{x})\leq\mathbf{0}\}$ is nonempty, closed, and convex with a Slater point.
VI-B Main result
We adopt the quadratic OLF (21) and optimizer $\mathbf{u}(\mathbf{z},t)$ as in (15) to impose the chosen decay law.
Theorem VI.2 (Strongly monotone games).
Under Assumption VI.1, the v-GNE $\mathbf{z}^{\star}$ solving (60a)-(60b) exists and is unique. For the closed loop dynamics in (15), the trajectory satisfies $S(\mathbf{z}(t))\to\mathbf{0}$ with the corresponding guarantee, as presented in (16).
Note that Assumption III.2 holds automatically in this setting. Furthermore, as discussed in Remark IV.7, radial unboundedness of the Lyapunov function is not required, since the primal–dual dynamics are expressed in feedback form and inherently guarantee bounded trajectories and convergence to the v-GNE.
Proof.
Existence and uniqueness of the variational equilibrium follow from strong monotonicity of $\mathcal{G}$ and convexity of the shared feasible set, as established by variational inequality theory [facchinei_generalized_2010, Sec. 3.2]. By construction of $\mathbf{u}(\mathbf{z},t)$, the OLF $V(\mathbf{z})$ satisfies the chosen decay law, hence $V(\mathbf{z}(t))\!\to\!0$ with the corresponding temporal guarantee, and thus $\mathbf{S}(\mathbf{z}(t))\!\to\!\mathbf{0}$.
To verify automatic nonsingularity, note that $V=\tfrac{1}{2}\|\mathbf{S}\|^{2}$ implies $\nabla V(\mathbf{z})=\nabla\mathbf{S}(\mathbf{z})^{\top}\mathbf{S}(\mathbf{z})$. For differentiable $\mathbf{S}$ induced by $\mathcal{G}$, strong monotonicity yields
$\frac{1}{2}\!\left(\nabla\mathbf{S}(\mathbf{z})+\nabla\mathbf{S}(\mathbf{z})^{\top}\right)\!\succeq m\,\mathbf{I}$ (65)
on the primal block corresponding to $\mathbf{x}$. Consequently,
$\mathbf{S}(\mathbf{z})^{\top}\nabla V(\mathbf{z})$ $=\mathbf{S}(\mathbf{z})^{\top}\nabla\mathbf{S}(\mathbf{z})\,\mathbf{S}(\mathbf{z})\;\geq\;m\,\|\mathbf{S}(\mathbf{z})\|^{2}.$ (66)
If $\mathbf{S}(\mathbf{z})\!\neq\!\mathbf{0}$, then the right-hand side of (66) is strictly positive, precluding $\nabla V(\mathbf{z})=\mathbf{0}$. Hence, Assumption III.2 holds automatically. ∎
VI-C Numerical Example: Cournot GNE with Shared Constraints
We illustrate Thm. VI.2 with a Cournot competition game [martinez-piazuelo_payoff_2022]. Consider $N=4$ firms and $M=2$ markets. Each firm selects a production vector $\mathbf{x}_{k}\in\mathbb{R}^{2}_{\geq 0}$, and the aggregate supply is $C\mathbf{x}$. Market prices follow a linear inverse demand
$\mathbf{J}(C\mathbf{x})=\bar{\mathbf{J}}-D\,C\mathbf{x},\qquad\bar{\mathbf{J}}=\begin{bmatrix}10\\ 8\end{bmatrix},\;D=\operatorname{diag}(1,1),$ (67)
and each firm incurs a quadratic production cost $Q_{k}(\mathbf{x}_{k})=\tfrac{1}{2}\|\mathbf{x}_{k}\|^{2}$. The shared constraints consist of one linear equality (a target supply in the first market) and market capacities,
$\mathbf{e}_{1}^{\top}C\mathbf{x}=12,\qquad C\mathbf{x}\preceq\begin{bmatrix}20\\ 15\end{bmatrix}$. (68)
The simulations in Fig. 3 confirm that the Cournot game behaves in line with the theoretical predictions of Theorem VI.2. The exponential law converges asymptotically, whereas the FT and FxT dynamics drive the stationarity vector to zero in bounded time. The PT design enforces convergence exactly at the user-specified horizon. The CPU-time results in Table III show that all four designs achieve convergence with comparable computational effort, with the PT case being particularly efficient in this instance. In addition, the trajectories of the multipliers and aggregate market quantities verify that both the linear equality and convex inequality constraints remain satisfied along the solution path, thereby validating the applicability of the OLF-based dynamics in this game-theoretic setting.
Figure 3: Decay of the optimality Lyapunov function $V(z(t))=\tfrac{1}{2}\|S(z(t))\|^{2}$ for the Cournot game under the four convergence laws. Exp converges asymptotically, FT and FxT converge in bounded time, and PT enforces convergence exactly at the prescribed horizon.
TABLE III: Wall-clock CPU time (ms) for the Cournot GNE.
Exp FT FxT PT
234 88 93 48
It is also worth noting that the observation of Remark IV.8 carries over unchanged here: since the Lyapunov function is built directly from the stacked stationarity vector, driving $V(\mathbf{z})$ below a given tolerance automatically certifies that each optimality component (stationarity, feasibility, and complementarity) is satisfied to the same numerical accuracy. This single scalar measure thus provides a unified guarantee for convergence to the v-GNE.
The Cournot game thus illustrates how the proposed control-centric dynamics extend beyond classical convex optimization to structured multi-agent equilibria, while retaining rigorous guarantees on convergence and constraint satisfaction.
VII Discussion About Discrete-Time Implementation
This section discusses the discrete-time realization of our continuous-time designs. Several recent directions inform the discrete-time design of continuous-time optimization dynamics. Discrete-gradient (DG) methods construct a step that exactly decreases a given Lyapunov/energy function, thereby preserving stability by construction [hernandez-solano_numerical_2023]. Related effective discretization schemes based on the Runge-Kutta (RK) family can preserve the Lyapunov decrease while reducing computational cost. In particular, reduced RK (RRK) methods simplify the standard RK update by removing numerically insignificant terms, maintaining the same stability properties with fewer operations [guedes_preservation_2025]. For homogeneous systems, Lyapunov-based discretizations can preserve exponential, FT, and FxT convergence rates and the Lyapunov function across the continuous-to-discrete transition [sanchez_lyapunov-based_2021]. More generally, a unified framework for translating Lyapunov arguments into discrete-time updates was proposed in [ushiyama_unified_2023], ensuring that discrete algorithms inherit the same Lyapunov guarantees as their continuous-time counterparts. FT and FxT guarantees for convex optimization, including proximal and minimax variants, have been established in continuous-time and analyzed under forward-Euler discretization (with small step sizes) [garg_fixed-time_2023]. These results suggest that discrete implementations should be guided by Lyapunov inequalities rather than tied to a specific numerical scheme.
VIII Conclusions
This paper has presented a control-centric framework for the systematic design of continuous-time optimization algorithms. The framework is founded on the concept of optimization Lyapunov functions (OLFs), constructed by selecting a Lyapunov candidate that encodes the optimality conditions and pairing it with a law selector that specifies the desired convergence rate. This formulation enables the synthesis of feedback dynamics with guaranteed exponential, FT, FxT, or PT convergence. Within this setting, three realizations were developed: the Hessian-gradient, Newton, and gradient dynamics, which differ in structure and information requirements but share the same Lyapunov-based synthesis principle. The Hessian-gradient dynamics leverage second-order information through the Hessian while avoiding explicit matrix inversion, providing a practical compromise between first- and second-order methods. The Newton dynamics employ the full Hessian feedback to achieve fast convergence, whereas the gradient dynamics represent the simplest realization based solely on first-order information. The framework was further applied to several problem classes: for unconstrained problems, it yielded guaranteed convergence for all dynamic realizations; for constrained optimization, Lyapunov-consistent KKT dynamics were constructed using smoothed stationarity mappings and shown to converge under all timing laws; for minimax problems, the same methodology extended convergence guarantees beyond the asymptotic and exponential cases available in prior work; and for GNE seeking, shared constraints were incorporated into the OLF design to ensure convergence to equilibrium points.
Beyond these specific results, the framework highlights a systematic path for embedding desired convergence profiles directly into optimization dynamics. This perspective provides design tools that complement existing algorithmic approaches and may inform future developments in intelligent optimization and cyber-physical decision-making systems.
Appendix A Proofs
A-A Proof of Corollary IV.6
Proof.
Strong convexity of $J$ implies the objective $J$ is strictly convex; affine equality constraints and convex inequalities define a convex feasible set. Under Slater, KKT conditions are necessary and sufficient for optimality; under LICQ and strong convexity, the primal solution $\mathbf{x}^{\star}$ is unique. Moreover, LICQ ensures uniqueness of the Lagrange multipliers $(\boldsymbol{\lambda}^{\star},\boldsymbol{\mu}^{\star})$ associated with $\mathbf{x}^{\star}$ [beck_introduction_2014, Thm. 11.12]. Hence, the KKT point is unique.
Let $V(\mathbf{z}):=\tfrac{1}{2}\|\mathbf{S}(\mathbf{z})\|^{2}$. Then $\nabla V(\mathbf{z})=\nabla\mathbf{S}(\mathbf{z})^{\top}\mathbf{S}(\mathbf{z})$. Furthermore, let $\mathbf{z}=(\mathbf{x},\boldsymbol{\lambda},\boldsymbol{\mu})$ and suppose $\nabla V(\mathbf{z})=\mathbf{0}$. Writing the stationarity of $V$ with respect to $(\boldsymbol{\mu},\boldsymbol{\lambda},\mathbf{x})$ yields:
$\mathbf{0}$ $=\frac{\partial V}{\partial\boldsymbol{\mu}}=\nabla\mathbf{h}(\mathbf{x})\,\nabla_{\mathbf{x}}L(\mathbf{x},\boldsymbol{\lambda},\boldsymbol{\mu})$
$\qquad\Rightarrow\quad\mathbf{A}\,\nabla_{\mathbf{x}}L(\mathbf{x},\boldsymbol{\lambda},\boldsymbol{\mu})=\mathbf{0},$ (69a)
$\mathbf{0}$ $=\frac{\partial V}{\partial\boldsymbol{\lambda}}=\nabla\mathbf{g}(\mathbf{x})\,\nabla_{\mathbf{x}}L(\mathbf{x},\boldsymbol{\lambda},\boldsymbol{\mu})$
$\qquad\qquad+\mathbf{D}_{a}(\boldsymbol{\lambda},\mathbf{g}(\mathbf{x}))\,\boldsymbol{\phi}_{\varepsilon}(\boldsymbol{\lambda},\mathbf{g}(\mathbf{x})),$ (69b)
$\mathbf{0}$ $=\frac{\partial V}{\partial\mathbf{x}}=\nabla^{2}_{\mathbf{x}\mathbf{x}}L(\mathbf{x},\boldsymbol{\lambda},\boldsymbol{\mu})\,\nabla_{\mathbf{x}}L(\mathbf{x},\boldsymbol{\lambda},\boldsymbol{\mu})+\nabla\mathbf{h}(\mathbf{x})^{\top}\mathbf{h}(\mathbf{x})$
$\qquad\qquad+\nabla\mathbf{g}(\mathbf{x})^{\top}\mathbf{D}_{b}(\boldsymbol{\lambda},\mathbf{g}(\mathbf{x}))\,\boldsymbol{\phi}_{\varepsilon}(\boldsymbol{\lambda},\mathbf{g}(\mathbf{x})),$ (69c)
where $\mathbf{D}_{a}=\mathrm{diag}(\alpha_{i})$ and $\mathbf{D}_{b}=\mathrm{diag}(\beta_{i})$ with (componentwise, for each inequality $i$)
$\alpha_{i}=\frac{\partial\phi_{\varepsilon}}{\partial\lambda_{i}}(\lambda_{i},g_{i})=\frac{\lambda_{i}}{\sqrt{\lambda_{i}^{2}+g_{i}^{2}+\varepsilon^{2}}}-1,$ (70a)
$\beta_{i}=\frac{\partial\phi_{\varepsilon}}{\partial g_{i}}(\lambda_{i},g_{i})=\frac{g_{i}}{\sqrt{\lambda_{i}^{2}+g_{i}^{2}+\varepsilon^{2}}}-1.$ (70b)
Since $\varepsilon>0$, we have $\alpha_{i},\beta_{i}\in(-1,0)$ for all $(\lambda_{i},g_{i})\in\mathbb{R}^{2}$, i.e., $\mathbf{D}_{a}\prec\mathbf{0}$ and $\mathbf{D}_{b}\prec\mathbf{0}$ are strictly negative diagonal.
Premultiplying (69c) by $\nabla_{\mathbf{x}}L$ and using (69a) gives
$\underbrace{\nabla_{\mathbf{x}}L^{\top}\,\nabla^{2}_{\mathbf{x}\mathbf{x}}L\,\nabla_{\mathbf{x}}L}_{\geq m\|\nabla_{\mathbf{x}}L\|^{2}}$ $\;+\;\nabla_{\mathbf{x}}L^{\top}\nabla\mathbf{g}^{\top}\mathbf{D}_{b}\,\boldsymbol{\phi}_{\varepsilon}$
$\quad+\;\underbrace{\nabla_{\mathbf{x}}L^{\top}\nabla\mathbf{h}^{\top}\mathbf{h}}_{=~(\mathbf{A}\nabla_{\mathbf{x}}L)^{\top}\mathbf{h}~=~0}~=~0.$ (71)
Strong convexity of $J$ implies $\nabla^{2}J(\mathbf{x})\succeq m\mathbf{I}$. Since each $g_{i}$ is convex, $\nabla^{2}g_{i}(\mathbf{x})\succeq\mathbf{0}$; and $\nabla^{2}h_{j}(\mathbf{x})=\mathbf{0}$ because $\mathbf{h}$ is affine. Therefore
$\nabla^{2}_{\mathbf{x}\mathbf{x}}L(\mathbf{x},\boldsymbol{\lambda},\boldsymbol{\mu})=\nabla^{2}J(\mathbf{x})+\sum_{i=1}^{p}\lambda_{i}\nabla^{2}g_{i}(\mathbf{x})~\succeq~m\mathbf{I},$ (72)
and the underbraced term in (A-A) satisfies $\nabla_{\mathbf{x}}L^{\top}\nabla^{2}_{\mathbf{x}\mathbf{x}}L\,\nabla_{\mathbf{x}}L\geq m\|\nabla_{\mathbf{x}}L\|^{2}$.
Next, from (69b), we have
$\nabla\mathbf{g}(\mathbf{x})\,\nabla_{\mathbf{x}}L(\mathbf{x},\boldsymbol{\lambda},\boldsymbol{\mu})=-\mathbf{D}_{a}\,\boldsymbol{\phi}_{\varepsilon}.$ (73)
Thus
$\nabla_{\mathbf{x}}L^{\top}$ $\nabla\mathbf{g}^{\top}\mathbf{D}_{b}\,\boldsymbol{\phi}_{\varepsilon}=\big(\nabla\mathbf{g}\,\nabla_{\mathbf{x}}L\big)^{\top}\mathbf{D}_{b}\,\boldsymbol{\phi}_{\varepsilon}$
$=\big(-\mathbf{D}_{a}\,\boldsymbol{\phi}_{\varepsilon}\big)^{\top}\mathbf{D}_{b}\,\boldsymbol{\phi}_{\varepsilon}=-\,\boldsymbol{\phi}_{\varepsilon}^{\top}\,\mathbf{D}_{a}\,\mathbf{D}_{b}\,\boldsymbol{\phi}_{\varepsilon}.$ (74)
Since $\mathbf{D}_{a}\prec\mathbf{0}$ and $\mathbf{D}_{b}\prec\mathbf{0}$ are diagonal, their product $\mathbf{D}_{a}\mathbf{D}_{b}$ is diagonal positive definite (each diagonal entry is the product of two strict negatives). Hence $\boldsymbol{\phi}_{\varepsilon}^{\top}\mathbf{D}_{a}\mathbf{D}_{b}\boldsymbol{\phi}_{\varepsilon}\leq 0$ with equality if and only if $\boldsymbol{\phi}_{\varepsilon}=\mathbf{0}$.
Therefore, from (A-A) we deduce
$0=\nabla_{\mathbf{x}}L^{\top}\nabla^{2}_{\mathbf{x}\mathbf{x}}L\,\nabla_{\mathbf{x}}L-\boldsymbol{\phi}_{\varepsilon}^{\top}\mathbf{D}_{a}\mathbf{D}_{b}\boldsymbol{\phi}_{\varepsilon}\geq m\|\nabla_{\mathbf{x}}L\|^{2},$ (75)
which implies $\nabla_{\mathbf{x}}L(\mathbf{x},\boldsymbol{\lambda},\boldsymbol{\mu})=\mathbf{0}$. Plugging back into (69b)-(69c) yields $\boldsymbol{\phi}_{\varepsilon}=\mathbf{0}$ and, with $\nabla_{\mathbf{x}}L=\mathbf{0}$ in (69c), $\mathbf{A}^{\top}\mathbf{h}(\mathbf{x})=\mathbf{0}$. Since $\mathbf{A}$ has full row rank, $\mathbf{h}(\mathbf{x})=\mathbf{0}$. Hence $\mathbf{S}(\mathbf{z})=\mathbf{0}$.
We have shown that $\nabla V(\mathbf{z})=\mathbf{0}$ implies $\mathbf{S}(\mathbf{z})=\mathbf{0}$. Therefore, $\nabla V(\mathbf{z})=\nabla\mathbf{S}(\mathbf{z})^{\top}\mathbf{S}(\mathbf{z})\neq\mathbf{0}$ whenever $\mathbf{S}(\mathbf{z})\neq\mathbf{0}$, i.e., Assumption III.2 holds automatically.∎